A drug is being delivered into a patient’s arm at a rate of 12.5 mL min–1. The drug is being delivered from a syringe through a 5 cm long needle with a radius of 0.75 mm. If the patient’s blood pressure is 120 kPa, what must the pressure in the syringe be? (ηdrug = 4.90×10-5 Pa s


1
Expert's answer
2021-03-21T11:25:39-0400
"P_g=\frac{\Delta V}{\Delta t}\frac{8L\eta}{\pi r^4}\\
P_g=\frac{0.0125}{60}\frac{8(0.05)(4.9\cdot10^{-5})}{\pi(0.75\cdot10^{-3})^4}=4108\ Pa"